8. A parallelepiped is formed by the vectors a,1,0, b 0,1, 1 and c1,1 e the volume of the parallelepiped. etermin 9. Show that the four points A(1,3, 2), B(3,-1,6), C(5, 2,0) and D(3,6,-4) are copla- nar
Answer #1

Solution : ( 8 )

\mathrm{Volume\:of\:parallelepiped\:V\:=}\:\left | a\cdot (b\times c) \right |

=\:<1,1,0>\cdot \begin{vmatrix} i & j& k\\ 0 & 1& 1\\ 1& 1& 1 \end{vmatrix}

=<1,1,0>\cdot \begin{vmatrix} 1 & 1\\ 1& 1 \end{vmatrix}i-\begin{vmatrix} 0 & 1\\ 1& 1 \end{vmatrix}j+\begin{vmatrix} 0 & 1\\ 1& 1 \end{vmatrix}k

=<1,1,0>\cdot <0,\:-1,\:-1>

=\left | 0+(-1)+(0) \right |

=\left | -1 \right |

=1

\mathrm{Hence,\:\:the\:volume\:of\:the\:parallelepiped\:is\:1.}

-----------------------------------------------------------------------------------------------------------------------------